OpenStudy (anonymous):

use the quadratic formula to solve : 2x^2-5x=1

OpenStudy (anonymous):

\[5\pm \sqrt{17}/4\]

OpenStudy (anonymous):

\[-b \pm \sqrt{b^2-4ac}/2a is your general form of \] is your general form of quadratic eq

OpenStudy (anonymous):

is that... \[5+\sqrt{17}/4 \] as in 5+or- (sqrt) 17 over 4?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

is the answer satisfactory ?

jhonyy9 (jhonyy9):

2x^2 -5x =1 2x^2 -5x -1 =0 x_1,2 = (5+/- sqrt(25+8)/4 = (5+/- sqrt33)/4

OpenStudy (anonymous):

yes it is